State the relation between acceleration and momentum?


1
Expert's answer
2021-06-07T09:32:36-0400

From the second Newton's law,

"\displaystyle F = ma = \frac{dp}{dt}"

So,

"\displaystyle a = \frac{1}{m} \frac{dp}{dt}"